Max Wavelength of Series for Hydrogen

What is the maximum wavelength seen in each of the Lyman, Balmer, Paschen, and Brackett, and Pfund series for hydrogen? (122, 656, 1880, 4050, 7470 nm)?
Answer #1

max wavelength happens with transition from n+1 to n
we use 1/λ=R(1/nf^2-1/n^2)
Lyman: n'=1

1/λ=1.097E7(1/1^2-1/2^2) so λ=1.22E-7 m

Balmer: n'=2

1/λ=1.097E7(1/2^2-1/3^2) so λ=6.56E-7 m

Paschen: n'=3

1/λ=1.097E7(1/3^2-1/4^2) so λ=1.88E-6 m

Bracket: n'=4

1/λ=1.097E7(1/4^2-1/5^2) so λ=4.05E-6 m

Pfund: n'=5

1/λ=1.097E7(1/5^2-1/6^2) so λ=7.46E-6 m

answered by: Brigid